### v1

@@ 5.3 Lower Bound for $L_{\infty}$ Symmetric
$(\hat{p}_{i,u,j,v,b})_{\ell}$ =$=\begin{cases}u&\text{if }\ell\in\{i,i+k\},\\ v&\text{if }\ell\in\{j,j+k\},\\ b&\text{otherwise,}\end{cases}\qquad(b_{i,j})_{\ell}=\begin{cases}10N&\text{if}\ell\in\{i,j\},\\ -10N&\text{if }\ell\in\{i+k,j+k\},\\ 0&\text{otherwise,}\end{cases}$
$(q_{i})_{\ell}$ =$=\begin{cases}10N&\text{if }\ell\in\{i,i+k\},\\ 0&\text{otherwise.}\end{cases}$
@@
$B^{\prime}_{i,j}$$\coloneqq\{(0,\ldots,0)\},$ $R^{\prime}_{i,s}$$\coloneqq\{-q_{i}\}.$
@@
Since the last coordinate of any point in $\hat{R}_{i,j}$ (or $\hat{R}^{\prime}_{i,j}$) is 0, its distance to $(0,\ldots,0,10N)+\tau$ is at least $10N+\tau_{d}$, with equality if $\tau\in[N]^{d}$. Similarly, its distance to $(0,\ldots,0,-10N)+\tau$ is at least $10N-\tau_{d}$, with equality if $\tau\in[N]^{d}$. Therefore, all copies of $(0,\ldots,0,10N)$ and $(0,\ldots,0,-10N)$ in total contribute a cost of at least $2(|E|-1)\cdot 10N$, with equality if $\tau\in[N]^{d}$, no matter what points in $\hat{B}_{i,j}$ (or $\hat{B}^{\prime}_{i,j}$, resp.) they are assigned to.
@@
We now apply the Gadget Combination Lemma to the gadgets $B_{i,s},R_{i,s},B^{\prime}_{i,s},R^{\prime}_{i,s}$ for $i\in[k]$ and $s\in[2(k-1)]$ and $\hat{B}_{i,j},\hat{R}_{i,j},\hat{B}^{\prime}_{i,j},\hat{R}^{\prime}_{i,j}$ for $1\leq i<j\leq k$. The $\mathrm{EMDuT}_{\infty}$ cost of the resulting point sets $B,R$ is the sum of the costs of the gadgets. Hence, cost can be read off from the above lemmas. Note that for any $i\in[k]$ the terms $|\tau_{i}-\tau_{i+k}|$ cancel, as there are $2(k-1)$ gadget pairs contributing a $+|\tau_{i}-\tau_{i+k}|$ (one gadget pair for each $s\in[2(k-1)]$), and $k-1$ gadget pairs contributing a $-2|\tau_{i}-\tau_{i+k}|$ (one gadget pair for each $j\in[k],j\neq i$). The cost thus simplifies to
@@
We obtain equality for any translation $\tau\in[N]^{d}$ with $(\tau_{i},\tau_{j})=(\tau_{i+k},\tau_{j+k})\in E$ for all $1\leq i<j\leq k$. Therefore, if $G$ has a $k$-Clique $v_{1},\ldots,v_{k}$, then $\tau\coloneqq(v_{1},\ldots,v_{k},v_{1},\ldots,v_{k},0)\in[N]^{d}$ has a total cost of $\Lambda$. On the other hand, if $G$ has no $k$-Clique, then for any $\tau\in\mathbb{R}^{d}$ there exist $1\leq i<j\leq k$ with $(\lfloor\tau_{i}\rceil,\lfloor\tau_{j}\rceil)\not\in E$ (as otherwise $(\lfloor\tau_{1}\rceil,\ldots,\lfloor\tau_{k}\rceil)$ would form a $k$-Clique). Then according to Lemma 5.7 at least one summand has an additional $+1$, so the total cost is at least $\Lambda+1$.
For any $\varepsilon<1/\Lambda$, a $(1+\varepsilon)$-approximation algorithm for $\mathrm{EMDuT}_{\infty}$ could distinguish cost at most $\Lambda$ and cost at least $\Lambda+1$, and thus would solve the $k$-Clique problem. Hence, if we would have a $(1+\varepsilon)$-approximation algorithm for $\mathrm{EMDuT}_{\infty}$ running in time $(n/\varepsilon)^{o(d)}$, then by setting $\varepsilon\coloneqq 0.9/\Lambda$ and observing $n=\mathcal{O}(N^{2}),1/\varepsilon=\mathcal{O}(\Lambda)=\mathcal{O}(N^{3})$, and $d=\mathcal{O}(k)$, we would obtain an algorithm for $k$-Clique running in time $(n/\varepsilon)^{o(d)}=\mathcal{O}(N^{5})^{o(k)}=N^{o(k)}$, which contradicts ETH by Theorem 5.2.
@@ 6 Algorithms in higher dimensions
Given two sets $R$ and $B$ of $n$ points in the plane, Eppstein et al. [23] show how to compute a translation $\tau^{*}$ minimizing $\mathrm{EMDuT}_{1}(B,R)$ with respect to the $L_{1}$-distance in $\mathcal{O}(n^{6}\log^{3}n)$ time. We observe that their result can be generalized to point sets in arbitrary dimension $d$, leading to an $\mathcal{O}(m^{d}n^{d+2}\log^{d+2}n)$ time algorithm.
Furthermore, we show that our approach can also be used to obtain an $\mathcal{O}(m^{d}n^{d+2}\log^{d+2}n)$ time algorithm for finding a translation that minimizes $\mathrm{EMDuT}_{\infty}(B,R)$, i.e. the Earth Mover’s Distance with respect to the $L_{\infty}$ distance. For point sets in $\mathbb{R}^{2}$, this immediately follows by “rotating the plane by $45^{\circ}$” and using the algorithm for $L_{1}$. For higher dimensions this trick is no longer immediately applicable. However, we show that our algorithm can also directly be applied to the $L_{\infty}$ distance, even for point sets in $\mathbb{R}^{d}$, with $d>2$.
@@
We first describe an algorithm to compute $\mathrm{EMD}_{p}(B,R)$ in $\mathbb{R}^{d}$. Note that we assume to work in the Real RAM model, hence we need a strongly-polynomial algorithm. Naively, one can achieve that in $\mathcal{O}(m^{2}n)$ time by computing the bipartite graph, and solving maximum weight matching in bipartite graph in strongly polynomial time by Edmonds and Karp [22]. Here, however, we can use the fact that points are in $\mathbb{R}^{d}$. To the best of our knowledge, the best algorithm in this setting is due to Vaidya [43]. However, he only considers the case when both point sets are in $\mathbb{R}^{2}$ and have size $n=m$ in $\mathbb{R}^{2}$. He shows that one can compute $\mathrm{EMD}_{p}(B,R)$ (with $p\in\{1,\infty\}$) in $\mathcal{O}(n^{2}\log^{3}n)$ time in this setting. Furthermore, he states (without proof) that for point sets in $\mathbb{R}^{d}$, that the running time increases by at most $\mathcal{O}(\log^{d}n)$. Next, we briefly sketch the algorithm and fill in the missing details for the higher-dimensional setting, to obtain the following result:
@@
Vaidya’s algorithm is a particular implementation of the Hungarian Method [36]. Hence, to apply it we introduce $n-m$ additional “dummy” points in $B$. We define the distance from a dummy point to any other point in $R$ to be zero. We stress that the dummy points are only present in the graph representation, and are not physically in $\mathbb{R}^{d}$. The main algorithm proceeds in phases, in each of which the current matching grows by one new pair. Hence, the algorithm completes after $n$ phases. In each phase, each point $q\in B\cup R$ is assigned a weight $w_{q}$ (the current value of the point in the dual LP-formulation of the problem). The algorithm then maintains a subset $R^{\prime}$ of red points (initially $R^{\prime}=R$), and a subset $B^{\prime}$ of blue points (initially, $B^{\prime}$ is the set of unmatched blue points), and repeatedly computes the bichromatic closest pair $(b^{*},r^{*})$ with respect to the weighted distance function $d(b,r)=\|br\|_{p}-w_{b}-w_{r}$. This pair is either added to the matching (if $r^{*}$ is also unmatched), or the weights of $b^{*}$ and $r^{*}$ are updated, $r^{*}$ is removed from $R^{\prime}$, and $b^{*}$ is added to $B^{\prime}$. Since there are only $n$ points in one set, a phase consists of at most $\mathcal{O}(n)$ such steps. Hence, given a dynamic data structure storing $R^{\prime}$ and $B^{\prime}$ that
@@
We describe a dynamic data structure for weighted nearest neighbor (NN) queries and then apply a recent result of Chan [15] to turn this into a dynamic bichromatic closest pair data structure.
Let $P^{\prime}$ be a set of weighted points in $\mathbb{R}^{d}$, let $q$ be a weighted point, and consider the positive orthant $A(q)=\{a\mid q_{i}\leq a_{i}\text{ for all }i\in\{1,\ldots,d\}\}\subset\mathbb{R}^{d}$ with respect to $q$ (i.e. all points dominating $q$). Let $w^{\prime}_{z}=-w_{p^{\prime}}+\sum_{i=1}^{d}p^{\prime}_{i}$ (see Figure 6.1). Now observe that the point $p^{*}\in P^{\prime}\cap A(q)$ minimizing $\|p^{\prime}q\|-w_{p^{\prime}}-w_{q}$ is the point in $A(q)$ with minimum $w^{\prime}_{p^{\prime}}$ value.
@@
So, we store the points in $P^{\prime}$ in a $d$-dimension range tree [39], in which every subtree is annotated with the point with the minimum $w^{\prime}_{p}$ value among its descendants. We can query the range tree for the nearest point among $P^{\prime}\cap A(q)$ with respect to the weighted $L_{1}$ distance in $\mathcal{O}(\log^{d}n)$ time. By implementing the trees using generalized balanced trees [5] we can support insertions and deletions in amortized $\mathcal{O}(\log^{d}n)$ time as well. We use an analogous approach for the other $12^{d}-$ 1 orthant around $q$, thus allowing us to answer weighted NN queries with respect to the $L_{1}$ distance in the same time as above. We can handle the (weighted) $L_{\infty}$ distance analogously by defining appropriate “orthants” around $q$. If $P^{\prime}$ contains any dummy points, we store them separately in a binary search tree, ordered by weight. The minimum weight dummy point is an additional candidate nearest neighbor for $q$, and we return the overall closest point to $q$. When $q$ itself would be a dummy point (and thus has no real location), we simply obtain the minimum weight point among $P^{\prime}$ as the nearest point. Hence, it follows we can build the structure in $P_{0}(n)=O(n\log^{d}n)$ time, query the nearest neighbor in $Q_{0}(n)=O(\log^{d}n)$ time, and delete (and insert) points in $D_{0}(n)=O(\log^{d}n)$ time.
Chan [15] shows how to turn a dynamic nearest neighbor searching data structure with preprocessing time $P_{0}(n)$, query time $Q_{0}(n)$, and deletion time $D_{0}(n)$ into a fully dynamic bichromatic closest pair data structure. In our particular case we obtain
@@
Plugging this in into Vaidia’s algorithm we thus obtain an $\mathcal{O}(n^{2}\log^{d+2}n)$ time algorithm to compute $\mathrm{EMD}(B,R)$. ∎
@@
The sets $B$ and $R$ are aligned in dimension $i$, or $i$-aligned for short, if there is a pair of points $b\in B,r\in R$ for which $b_{i}=r_{i}$. Eppstein et al. [23] show that for two point sets in $\mathbb{R}^{2}$, there exists an optimal translation $\tau^{*}$ that aligns $B$ and $R$ in both dimensions. They explicitly consider all $\mathcal{O}((nm)^{2})$ translations that both $1$-align and $2$-align $B+\tau$ and $R$. For each such a translation $\tau$, computing an optimal matching can then be done in $\mathcal{O}(n^{2}\log^{3}n)$ time [43], thus leading to an $\mathcal{O}(n^{4}m^{2}\log^{3}n)$ time algorithm. We now argue that we can generalize the above result to higher dimensions.
@@
The set $H$ contains $mnd$ hyperplanes, and thus $\mathcal{A}(H)$ contains $\mathcal{O}((mnd)^{d})=\mathcal{O}(m^{d}n^{d})$ vertices. Computing $\mathcal{A}(H)$ takes $\mathcal{O}(m^{d}n^{d})$ time [20, 21]. For each such a vertex (translation), we can compute an optimal matching in $\mathcal{O}(n^{2}\log^{d+2}n))$ time using the algorithm from Theorem 6.1. This thus yields an $\mathcal{O}(m^{d}n^{d+2}\log^{d+2}n)$ time algorithm in total. ∎
@@
In case of the $L_{\infty}$ distance, we use a similar approach as in Theorem 6.2. We prove that there is a set $H$ of $\mathcal{O}(mnd^{2})$ hyperplanes in $\mathbb{R}^{d}$, so that for any matching $\phi$, there is a minimum cost translation that is a vertex of the arrangement $\mathcal{A}(H)$. We can thus again compute such an optimal matching (and the translation) by trying all $\mathcal{O}(m^{d}n^{d})$ vertices. This yields the following result, thereby also establishing Theorem 1.3.
@@
which is again a piecewise linear function in $\tau$, as it is a sum of piecewise linear functions. Each individual function is now of the form $f_{b,r}=\max_{i=1}^{d}|b_{i}+\tau_{i}-r_{i}|$, where $b\in B$, and $r=\phi(b)$. For each such a function, there are now at most $\mathcal{O}(d)$ hyperplanes that subdivide $\mathbb{R}^{d}$ into regions in which $f_{b,r}$ is given by a single linear function. In particular, the $d$ hyperplanes given by $\tau_{i}-b_{i}+r_{i}=0$, for any $i\in\{1,\ldots,d\}$, and $\mathcal{O}(d^{2})$ hyperplanes that we get from solving $|b_{i}+\tau_{i}-r_{i}|=|b_{j}+\tau_{j}-r_{j}|$ for each $j\neq i$ (for example $\tau_{i}-\tau_{j}+b_{i}-b_{j}-r_{i}+r_{j}=0$). Let $H$ denote the resulting set of $\mathcal{O}(mnd^{2})$ hyperplanes over all points $b\in B$, $r\in R$, and all $i\in\{1,\ldots,d\}$. It again follows that for any matching $\phi$, there is a vertex of $\mathcal{A}(H)$ that corresponds to a translation $\tau_{\phi}$ that minimizes $\mathcal{D}_{B,R,\infty}(\phi,\tau)$. Since this holds for an optimal matching $\phi^{*}$, we can thus compute $\mathrm{EMDuT}(B,R)$ in $\mathcal{O}(m^{d}n^{d+2}\log^{d+2}n)$ time. ∎

### v2

@@ 5.3 Lower Bound for $L_{\infty}$ Symmetric
$(\hat{p}_{i,u,j,v,b})_{\ell}$ $=\begin{cases}u&\text{if }\ell\in\{i,i+k\},\\ v&\text{if }\ell\in\{j,j+k\},\\ b&\text{otherwise,}\end{cases}\qquad(b_{i,j})_{\ell}=\begin{cases}10N&\text{if }\ell\in\{i,j\},\\ -10N&\text{if }\ell\in\{i+k,j+k\},\\ 0&\text{otherwise,}\end{cases}$
$(q_{i})_{\ell}$ $=\begin{cases}10N&\text{if }\ell\in\{i,i+k\},\\ 0&\text{otherwise.}\end{cases}$
@@
$B^{\prime}_{i,s}$ $\coloneqq\{(0,\ldots,0)\},$ $R^{\prime}_{i,s}$ $\coloneqq\{-q_{i}\}.$
@@
Since the last coordinate of any point in $\hat{R}_{i,j}$ (or $\hat{R}^{\prime}_{i,j}$) is 0, its distance to $(0,\ldots,0,10N)+\tau$ is at least $10N+\tau_{d}$, with equality if $\tau\in[N]^{d}$. Similarly, its distance to $(0,\ldots,0,-10N)+\tau$ is at least $10N-\tau_{d}$, with equality if $\tau\in[N]^{d}$. Therefore, all copies of $(0,\ldots,0,10N)$ and $(0,\ldots,0,-10N)$ in total contribute a cost of at least $2(|E|-1)\cdot 10N$, with equality if $\tau\in[N]^{d}$, no matter what points in $\hat{R}_{i,j}$ (or $\hat{R}^{\prime}_{i,j}$, resp.) they are assigned to.
@@
We now apply the Gadget Combination Lemma to the gadgets $B_{i,s},R_{i,s},B^{\prime}_{i,s},R^{\prime}_{i,s}$ for $i\in[k]$ and $s\in[2(k-1)]$ and $\hat{B}_{i,j},\hat{R}_{i,j},\hat{B}^{\prime}_{i,j},\hat{R}^{\prime}_{i,j}$ for $1\leq i<j\leq k$. The value of $\mathrm{EMDuT}_{\infty}$ of the resulting point sets $B,R$ is the sum of the costs of the gadgets. Hence, the cost can be read off from the above lemmas. Note that for any $i\in[k]$ the terms $|\tau_{i}-\tau_{i+k}|$ cancel, as there are 2(k1)2(k-1) gadget pairs contributing $|\tau_{i}-\tau_{i+k}|$ (one gadget pair for each $s\in[2(k-1)]$), and k1k-1 gadget pairs contributing $2|\tau_{i}-\tau_{i+k}|$ (one gadget pair for each $j\in[k],j\neq i$). The cost thus simplifies to
@@
We obtain equality for any translation $\tau\in[N]^{d}$ with $(\tau_{i},\tau_{j})=(\tau_{i+k},\tau_{j+k})\in E$ for all $1\leq i<j\leq k$. Therefore, if $G$ has a $k$-Clique $v_{1},\ldots,v_{k}$, then $\tau\coloneqq(v_{1},\ldots,v_{k},v_{1},\ldots,v_{k},0)\in[N]^{d}$ has a total cost of $\Lambda$. On the other hand, if $G$ has no $k$-Clique, then for any $\tau\in\mathbb{R}^{d}$ there exist $1\leq i<j\leq k$ with $(\lfloor\tau_{i}\rceil,\lfloor\tau_{j}\rceil)\not\in E$ (as otherwise $(\lfloor\tau_{1}\rceil,\ldots,\lfloor\tau_{k}\rceil)$ would form a $k$-Clique). Then according to Lemma 5.7 at least one summand has an additional +1+1, so the total cost is at least $\Lambda+1$.
For any $\varepsilon<1/\Lambda$, a $(1+\varepsilon)$-approximation algorithm for $\mathrm{EMDuT}_{\infty}$ could distinguish cost at most $\Lambda$ and cost at least $\Lambda+1$, and thus would solve the $k$-Clique problem. Hence, if we had a $(1+\varepsilon)$-approximation algorithm for $\mathrm{EMDuT}_{\infty}$ running in time $(n/\varepsilon)^{o(d)}$, then by setting $\varepsilon\coloneqq 0.9/\Lambda$ and observing $n=\mathcal{O}(N^{2}),1/\varepsilon=\mathcal{O}(\Lambda)=\mathcal{O}(N^{3})$, and $d=\mathcal{O}(k)$, we would obtain an algorithm for $k$-Clique running in time $(n/\varepsilon)^{o(d)}=\mathcal{O}(N^{5})^{o(k)}=N^{o(k)}$, which contradicts ETH by Theorem 5.2.
@@ 6 Algorithms in Higher Dimensions
Given two sets $R$ and $B$ of $n$ points in the plane, Eppstein et al. [eppstein15improv_grid_map_layout_point_set_match] show how to compute a translation $\tau^{*}$ minimizing $\mathrm{EMDuT}_{1}(B,R)$ with respect to the $L_{1}$-distance in $\mathcal{O}(n^{6}\log^{3}n)$ time. We observe that their result can be generalized to point sets in arbitrary dimension $d$, leading to an $\mathcal{O}(m^{d}n^{d+2}\log^{d+2}n)$ time algorithm.
Furthermore, we show that our approach can be used to obtain an $\mathcal{O}(m^{d}n^{d+2}\log^{d+2}n)$ time algorithm for finding a translation that minimizes $\mathrm{EMDuT}_{\infty}(B,R)$, i.e. the Earth Mover’s Distance with respect to the $L_{\infty}$ distance. For point sets in $\mathbb{R}^{2}$, this immediately follows by “rotating the plane by $45^{\circ}$” and using the algorithm for $L_{1}$. For higher dimensions this trick is no longer immediately applicable. However, we show that our algorithm can also directly be applied to the $L_{\infty}$ distance, even for point sets in $\mathbb{R}^{d}$ with $d>2$.
@@
We first describe an algorithm to compute $\mathrm{EMD}_{p}(B,R)$ in $\mathbb{R}^{d}$. Note that we assume to work in the Real RAM model, hence we need a strongly-polynomial algorithm. Naively, one can achieve that in $\mathcal{O}(m^{2}n)$ time by computing the bipartite graph, and solving maximum weight matching in bipartite graph in strongly polynomial time by Edmonds and Karp [edmonds1972theoretical]. Here, however, we can use the fact that points are in $\mathbb{R}^{d}$. To the best of our knowledge, the best algorithm in this setting is due to Vaidya [vaidya89geomet_helps_match]. However, he only considers the case when both point sets are in $\mathbb{R}^{2}$ and have size $n=m$ in $\mathbb{R}^{2}$. He shows that one can compute $\mathrm{EMD}_{p}(B,R)$ (with $p\in\{1,\infty\}$) in $\mathcal{O}(n^{2}\log^{3}n)$ time in this setting. Furthermore, he states (without proof) that for point sets in $\mathbb{R}^{d}$, that the running time increases by at most $\mathcal{O}(\log^{d}n)$. Next, we briefly sketch the algorithm and fill in the missing details for the higher-dimensional setting, to obtain the following result:
@@
Vaidya’s algorithm is a particular implementation of the Hungarian Method [kuhn1955hungarian]. Hence, to apply it we introduce nmn-m additional “dummy” points in $B$. We define the distance from a dummy point to any other point in $R$ to be zero. We stress that the dummy points are only present in the graph representation, and are not physically in $\mathbb{R}^{d}$. The main algorithm proceeds in phases, in each of which the current matching grows by one new pair. Hence, the algorithm completes after $n$ phases. In each phase, each point $q\in B\cup R$ is assigned a weight $w_{q}$ (the current value of the point in the dual LP-formulation of the problem). The algorithm then maintains a subset $R^{\prime}$ of red points (initially $R^{\prime}=R$), and a subset $B^{\prime}$ of blue points (initially, $B^{\prime}$ is the set of unmatched blue points), and repeatedly computes the bichromatic closest pair $(b^{*},r^{*})$ with respect to the weighted distance function $d(b,r)=\|b-r\|_{p}-w_{b}-w_{r}$. This pair is either added to the matching (if $r^{*}$ is also unmatched), or the weights of $b^{*}$ and $r^{*}$ are updated, $r^{*}$ is removed from $R^{\prime}$, and $b^{*}$ is added to $B^{\prime}$. Since there are only $n$ points in one set, a phase consists of at most $\mathcal{O}(n)$ such steps. Hence, given a dynamic data structure storing $R^{\prime}$ and $B^{\prime}$ that
@@
We describe a dynamic data structure for weighted nearest neighbor (NN) queries and then apply a recent result of Chan [chan20dynam_gener_closes_pair] to turn this into a dynamic bichromatic closest pair data structure.
Let $P^{\prime}$ be a set of weighted points in $\mathbb{R}^{d}$, let $q$ be a weighted point, and consider the positive orthant $A(q)=\{a\mid q_{i}\leq a_{i}\text{ for all }i\in\{1,\ldots,d\}\}\subset\mathbb{R}^{d}$ with respect to $q$ (i.e. all points dominating $q$). Let $w^{\prime}_{z}=-w_{p^{\prime}}+\sum_{i=1}^{d}p^{\prime}_{i}$ (see Figure 6.1). Now observe that the point $p^{*}\in P^{\prime}\cap A(q)$ minimizing $\|p^{\prime}-q\|-w_{p^{\prime}}-w_{q}$ is the point in $A(q)$ with minimum $w^{\prime}_{p^{\prime}}$ value.
@@
So, we store the points in $P^{\prime}$ in a $d$-dimension range tree [preparata2012computational], in which every subtree is annotated with the point with the minimum $w^{\prime}_{p}$ value among its descendants. We can query the range tree for the nearest point among $P^{\prime}\cap A(q)$ with respect to the weighted $L_{1}$ distance in $\mathcal{O}(\log^{d}n)$ time. By implementing the trees using generalized balanced trees [andersson99gener_balan_trees] we can support insertions and deletions in amortized $\mathcal{O}(\log^{d}n)$ time as well. We use an analogous approach for the other $2^{d}-1$ orthants around $q$, thus allowing us to answer weighted NN queries with respect to the $L_{1}$ distance in the same time as above. We can handle the (weighted) $L_{\infty}$ distance analogously by defining appropriate “orthants” around $q$. If $P^{\prime}$ contains any dummy points, we store them separately in a binary search tree, ordered by weight. The minimum weight dummy point is an additional candidate nearest neighbor for $q$, and we return the overall closest point to $q$. When $q$ itself would be a dummy point (and thus has no real location), we simply obtain the minimum weight point among $P^{\prime}$ as the nearest point. Hence, it follows we can build the structure in $P_{0}(n)=O(n\log^{d}n)$ time, query the nearest neighbor in $Q_{0}(n)=O(\log^{d}n)$ time, and delete (and insert) points in $D_{0}(n)=O(\log^{d}n)$ time.
Chan [chan20dynam_gener_closes_pair] shows how to turn a dynamic nearest neighbor searching data structure with preprocessing time $P_{0}(n)$, query time $Q_{0}(n)$, and deletion time $D_{0}(n)$ into a fully dynamic bichromatic closest pair data structure. In our particular case we obtain
@@
Plugging this into Vaidia’s algorithm we thus obtain an $\mathcal{O}(n^{2}\log^{d+2}n)$ time algorithm to compute $\mathrm{EMD}(B,R)$. ∎
@@
The sets $B$ and $R$ are aligned in dimension $i$, or $i$-aligned for short, if there is a pair of points $b\in B,r\in R$ for which $b_{i}=r_{i}$. Eppstein et al. [eppstein15improv_grid_map_layout_point_set_match] show that for two point sets in $\mathbb{R}^{2}$, there exists an optimal translation $\tau^{*}$ that aligns $B$ and $R$ in both dimensions. They explicitly consider all $\mathcal{O}((nm)^{2})$ translations that both $1$-align and $2$-align $B+\tau$ and $R$. For each such a translation $\tau$, computing an optimal matching can then be done in $\mathcal{O}(n^{2}\log^{3}n)$ time [vaidya89geomet_helps_match], thus leading to an $\mathcal{O}(n^{4}m^{2}\log^{3}n)$ time algorithm. We now argue that we can generalize the above result to higher dimensions.
@@
The set $H$ contains mndmnd hyperplanes, and thus $\mathcal{A}(H)$ contains $\mathcal{O}((mnd)^{d})=\mathcal{O}(m^{d}n^{d})$ vertices. Computing $\mathcal{A}(H)$ takes $\mathcal{O}(m^{d}n^{d})$ time [bkos-cgaa-08, edelsbrunner86const_arran_lines_hyper_applic]. For each such a vertex (translation), we can compute an optimal matching in $\mathcal{O}(n^{2}\log^{d+2}n))$ time using the algorithm from Theorem 6.1. This thus yields an $\mathcal{O}(m^{d}n^{d+2}\log^{d+2}n)$ time algorithm in total. ∎
@@
We present the following analog for $L_{\infty}$.
@@
which is again a piecewise linear function in $\tau$, as it is a sum of piecewise linear functions. Each individual function is now of the form $f_{b,r}=\max_{i=1}^{d}|b_{i}+\tau_{i}-r_{i}|$, where $b\in B$, and $r=\phi(b)$. For each such a function, there are now at most $\mathcal{O}(d^{2})$ hyperplanes that subdivide $\mathbb{R}^{d}$ into regions in which $f_{b,r}$ is given by a single linear function. In particular, the $d$ hyperplanes given by $\tau_{i}-b_{i}+r_{i}=0$, for any $i\in\{1,\ldots,d\}$, and $\mathcal{O}(d^{2})$ hyperplanes that we get from solving $|b_{i}+\tau_{i}-r_{i}|=|b_{j}+\tau_{j}-r_{j}|$ for each $j\neq i$ (for example $\tau_{i}-\tau_{j}+b_{i}-b_{j}-r_{i}+r_{j}=0$). Let $H$ denote the resulting set of $\mathcal{O}(mnd^{2})$ hyperplanes over all points $b\in B$, $r\in R$, and all $i\in\{1,\ldots,d\}$. It again follows that for any matching $\phi$, there is a vertex of $\mathcal{A}(H)$ that corresponds to a translation $\tau_{\phi}$ that minimizes $\mathcal{D}_{B,R,\infty}(\phi,\tau)$. Since this holds for an optimal matching $\phi^{*}$, we can thus compute $\mathrm{EMDuT}(B,R)$ in $\mathcal{O}(m^{d}n^{d+2}\log^{d+2}n)$ time. ∎